OpenStudy (trojanpoem):

if 6^ (-1/a) = 3^(1/b) = 2^ (1/g)

OpenStudy (trojanpoem):

Prove that 

OpenStudy (dan815):

is that the determinant

OpenStudy (trojanpoem):

Yeah , no expansion

ganeshie8 (ganeshie8):

what do we know about a,b,g are they integers

OpenStudy (trojanpoem):

6^ (-1/a) = 3^(1/b) = 2^ (1/g)

OpenStudy (dan815):